OpenStudy (anonymous):

If I have only 2 Sides of a triangle how do I find the 3rd?

OpenStudy (anonymous):

if you are looking for the length of the third side then you can use Pythagorean theorem

OpenStudy (anonymous):

law of cosines. \[c^2 = a^2 + b^2 - 2abcos(\theta)\] where \[\theta\]is the angle between those 2 sides. when the angle is 90, cos90 = 0 and it boils down to the pythagorean theorem.

OpenStudy (anonymous):

each triangle adds up to be 180 degrees so just add up the two that you have now and subtract it from 180 and you have your answer :)

OpenStudy (anonymous):

those are angle measurements, not side measurements

OpenStudy (anonymous):

if u have 2 of them whats the number of the 2 sides?